OpenStudy (kainui):

A silly question. I am getting lost somewhere in the math, how do I represent a power series as a fourier series?

OpenStudy (kainui):

\[\Large \sum_{n=0}^\infty a_nt^n = \sum_{m=-\infty}^\infty c_m e^{\frac{i 2\pi m }{T}t}\]